# Free Descriptions of Convex Sets

### v2

@@ 1.1.3 Invariant Conic Programs
Our framework yields structural results not only for descriptions of convex sets but also for sequences of optimization problems over them. Specifically, there is a literature showing that many sequences of invariant conic programs indexed by dimension can be solved in constant time. We use representation stability to unify and generalize these results. Such sequences of programs arise in several applications including extremal combinatorics [3, 4] and quantum information [5, 6], as we alluded to previously. Concretely, the programs that arise in certifying homomorphism density inequalities over graphs are invariant under symmetric groups of increasing sizes that relabel the vertices of the graphs involved. Similarly, optimizing traces of matrix polynomials yields programs that are invariant under the unitary or orthogonal groups of increasing sizes that conjugate the matrices involved. We now recall symmetry reductions of invariant programs and explain how constant-sized reductions arise from representation stability.
@@
Let $\{\mathbb{U}_{n}\}$ be a sequence of $\{\mathsf{G}_{n}\}$ representations and $\{\mathcal{K}_{n}\subseteq\mathbb{U}_{n}\}$ a sequence of convex cones. For $t\in\mathbb{N}$, we say that the sequence $\{\mathcal{K}_{n}^{\mathsf{G}_{n}}\subseteq\mathbb{U}_{n}^{\mathsf{G}_{n}}\}$ admits a constant-sized description for $n\geq t$ if there exists a single vector space $\mathbb{U}$ containing a cone $\mathcal{K}$, linear maps $T_{n}\colon\mathbb{U}\to\mathbb{U}_{n}^{\mathsf{G}_{n}}$, and subspaces $\mathbb{L}_{n}\subseteq\mathbb{U}$ such that $\mathcal{K}_{n}^{\mathsf{G}_{n}}=T_{n}(\mathcal{K}\cap\mathbb{L}_{n})$ for all $n\geq t$.
Note that Definition 1.10 does not require $\{\mathbb{U}_{n}\}$ to be a consistent sequence, nor in particular that $\mathbb{U}_{n}$ is embedded into $\mathbb{U}_{n+1}$. Proofs of constant-sized symmetry reductions in the literature have implicitly proceeded in a case-by-case manner by showing that the relevant cones, including symmetric PSD and relative entropy cones, have a constant-sized description in the sense of our Definition 1.10, see [25, 3, 26, 27]. In Section 5, we explain how these constant-sized descriptions can be generalized and derived systematically from an interplay between representation stability and the structure of the cones in question. To that end, we use a stronger form of representation stability known as uniform representation stability, which shows that the whole decomposition into irreducibles of the sequences of representations involved stabilize, rather than only their spaces of invariants; see Section 2.5. Our approach allows us to prove the following results.
Theorem 1.11 (Informal).
@@
We refer the reader to Theorem 5.1 for the formal statement and its proof. The sequences of classical Weyl groups are the groups of permutations $\{\mathsf{S}_{n}\}$, signed permutations $\{\mathsf{B}_{n}\}$, or even signed permutations $\{\mathsf{D}_{n}\}$. Here $\mathrm{Sym}^{2}_{+}(\mathbb{V})$ denotes the collection of nonnegative quadratic forms on $\mathbb{V}$. For example, Theorem 1.11 shows that the sections of the PSD cones $\{\mathbb{S}^{n}_{+}\}$ invariant under the usual action of the classical Weyl groups by conjugation admit constant-sized descriptions. Theorem 1.11 is used to derive constant-sized descriptions for cones of invariant sums-of-squares in Theorem 1.12 below. In the following result, we consider polynomials in $\binom{n}{k}$ variables identified with $k$-subsets of $n$ letters, and show that the cones of sums of squares of such polynomials modulo any sequence of ideals admit constant-sized descriptions. The proof of the following result is given in Section 5.1.
@@
$\mathcal{I}_{n}\subseteq\bigoplus_{d\geq 0}\mathrm{Sym}^{d}\left(\bigwedge\nolimits^{k}\mathbb{R}^{n}\right)\cong\mathbb{R}[x_{i_{1},\ldots,i_{k}}]_{1\leq i_{1}<\cdots<i_{k}\leq n}=:\mathbb{V}_{n}$
be $\mathsf{G}_{n}$-invariant ideals, and let $\mathbb{U}_{n}=\mathrm{Sym}^{\leq 2d}(\bigwedge^{k}\mathbb{R}^{n})/\mathcal{I}_{n}$. Consider the sums-of-squares cones $\mathrm{SOS}_{\mathbb{U}_{n}}=\{f\in\mathbb{U}_{n}:f\textrm{ is a sum of squares mod }\mathcal{I}_{n}\}$. Then the sequence $\{\mathrm{SOS}_{\mathbb{U}_{n}}^{\mathsf{G}_{n}}\}$ admits a constant-sized description for $n\geq 2kd$ if $\mathsf{G}_{n}=\mathsf{S}_{n}$ or $\mathsf{B}_{n}$, and for $n\geq 2kd+1$ if $\mathsf{G}_{n}=\mathsf{D}_{n}$.
If $k=1$, $\mathcal{I}_{n}=(0)$, and $\mathsf{G}_{n}=\mathsf{S}_{n}$ then we recover [25, Thms. 4.7, 4.10], and when $\mathsf{G}_{n}=\mathsf{B}_{n}$ or $\mathsf{D}_{n}$ we recover [26, Cor. 3.23]. If $k\geq 2$, if $\mathcal{I}_{n}=(x_{I}-x_{I}^{2})_{I\subseteq\binom{[n]}{k}}$ is the ideal generated by $x_{I}^{2}-x_{I}$ where $I$ ranges over all $k$-subsets of $[n]$, and if $\mathsf{G}_{n}=\mathsf{S}_{n}$, we recover [3, Thm. 2.4]. Theorem 1.12 generalizes all of these results to include any of the classical Weyl groups and any sequence of invariant ideals.
An application of Theorem 1.12 is obtaining constant-sized SDPs to certify graph homomorphism density inequalities [3, 4]. Many problems in extremal combinatorics can be recast as proving polynomial inequalities between homomorphism densities of graphs, which is the fraction of maps between the vertex sets of two graphs that define graph homomorphisms. A simple example is Mantel’s theorem, which states that the maximum number of edges in a triangle-free graph is $\lfloor n^{2}/4\rfloor$. Razborov proposed a method of certifying such inequalities using flag algebras [28], which were shown in [3, 4] to be sums-of-squares certificates of certain symmetric polynomial inequalities. Razborov’s flags are interpreted in [3, §3] as “free” spanning sets for spaces of symmetric polynomials. Formally, they are freely-described elements in the sense of Definition 1.4. Our framework shows that both the existence of such freely-described spanning sets and the resulting constant-sized SDPs are consequences of representation stability.
@@
1.1.4 Free Convex Regression
As our final contribution, we apply our framework to develop an algorithm for a free analog of the convex regression problem. In the classic setting of this problem, the objective is to fit a convex function $f\colon\mathbb{V}\to\mathbb{R}$ to a dataset $\{(x_{i},y_{i})\}_{i=1}^{D}$ of inputs $x_{i}\in\mathbb{V}$ and outputs $y_{i}\in\mathbb{R}$ such that $f(x_{i})\approx y_{i}$. In our free extension of this problem, we have a sequence $\{\mathbb{V}_{n}\}$ of vector spaces, usually of growing dimension, and data $\{(x_{i},y_{i})\in\mathbb{V}_{n_{i}}\oplus\mathbb{R}\}$ in finitely-many dimensions $n_{i}$. We then seek an infinite sequence of convex functions $\{f_{n}\colon\mathbb{V}_{n}\to\mathbb{R}\}$ that can be instantiated in any dimension, satisfying $f_{n_{i}}(x_{i})\approx y_{i}$ in the dimensions in which data is available.
We address this problem by leveraging the framework described above to fit a freely-described convex function to the given data. Concretely, we begin by endowing the vector spaces $\{\mathbb{V}_{n}\}$ containing the training data with the structure of a consistent sequence based on the symmetries of the problem at hand and the relations between problem instances in different dimensions. We then select description spaces $\{\mathbb{W}_{n}\},\{\mathbb{U}_{n}\}$ and cones $\{\mathcal{K}_{n}\subseteq\mathbb{U}_{n}\}$ with respect to which our convex functions are to be described as in (ConicSeq), with freely-described vectors and linear maps parametrizing the desired family of freely-described functions; richer description spaces generally yield more expressive families of convex functions, but fitting a function from such a family is more expensive and requires data in higher dimensions by Theorem 3.5.
@@ 1.1.4 Free Convex Regression
The $\ell_{\pi}$ norm of a vector is defined for vectors of any length, is invariant under signed permutations, remains unchanged under zero-padding of its input, and can only increase if we append any other entry to a given input. It is therefore both intersection and projection compatible with respect to the embeddings and projections of Example 1.5. The function (2) is well-defined for (positive-semidefinite) inputs of all sizes, is invariant under conjugation by orthogonal matrices, and its value is unchanged if we zero-pad its input. Thus, it is intersection-compatible on $\{\mathbb{S}^{n}\}$ with embeddings given by by zero-padding. Given evaluations of these two functions on low-dimensional inputs, we use our procedure to fit semidefinite approximations to them by choosing sequences of PSD cones for our descriptions, and fitting a freely-described and compatible sequence of convex functions. Thus, we ensure that our approximations are also group-invariant and satisfy the same compatibility properties as the underlying ground-truth functions (we detail our description spaces in Section 6.3)). Figure 1 shows the error in our semidefinite approximations for different dimensions when we search over all freely-described sets in our family when fitting to data, and when we search only over the smaller family of compatible sequences. We see that imposing compatibility ensures that the error increases gracefully when extending to higher dimensions.
@@
Obtaining semidefinite approximations for quantum information theoretic quantities such as (2) facilitates the use of semidefinite programming solvers in convex optimization problems involving quantum entropy [32, 33, 34, 35]. We mention that the authors of [36] derived semidefinite approximations of the quantum entropy and related functions from an analytic perspective. We show in Section 4.5 that their description is free but does not certify compatibility, in contrast to the description we obtain from data using the preceding procedure.
@@ 1.2 Related Work
There is a large literature as well as a systematic framework on extended formulations in which conic descriptions of a convex set in a fixed dimension are investigated; see [37] for a review. The goal in this body of work is to express ‘complicated’ convex sets in $\mathbb{R}^{d}$ (e.g., polyhedra with many facets) as linear images of affine sections of ‘simple’ convex cones in a space that is not much larger than $d$. This framework is also applied to study equivariant lifts of group-invariant convex sets, which are descriptions of the form (Conic) consisting of group-invariant cones, vectors, and linear maps, i.e., conic descriptions that make evident the group invariance of the convex set; see [38] and [37, §4.3]. These are precisely the type of descriptions we consider for each of the convex sets in our sequences. Moreover, while the literature on extended formulations is typically articulated in the setting of a convex set in a fixed dimension, many results in the area implicitly concern descriptions of a sequence $\{\mathcal{C}_{n}\}$ of convex sets and the complexity of these descriptions as a function of $n$ [39, 38, 40, 41]. Thus, there are several points of contact with the present paper. In fact, many (though not all) descriptions proposed in the literature on extended formulations can in fact be instantiated in any dimension, and are moreover free in our sense as we show in Section 4. More broadly, to the best of our knowledge, descriptions of convex sets that arise from a systematic consideration of relations between dimensions have not been studied previously. By bringing such considerations to the fore, the present effort elucidates the representation-theoretic phenomena and their interaction with the convex-geometric aspects underpinning convex sets that can be instantiated in any dimension.
@@
A broad research program pursued in several areas involves the study of “matrix” or noncommutative analogues of classical “scalar” or commutative objects. Examples include random matrix theory and free probability [42] in which matrix-valued random variables and their limits are the object of study as opposed to scalar-valued ones; and noncommutative algebraic geometry [13], in which polynomials in noncommuting variables and their evaluations on matrices are studied as opposed to standard polynomials in commuting variables that are evaluated on scalars [43, 44, 45, 46]. Applying this program to convex sets yields matrix-convex sets and free spectrahedra as in Example 1.2(b). We refer the reader to [13, 14] for surveys and [47, 48] for some applications. In analogy to the present paper, results in this area explicitly pertain to sequences of sets which are “freely-described”, in the sense that their description can be instantiated in any dimension. For example, free spectrahedra are sequences of sets described by a single linear matrix inequality, and free algebraic varieties are defined by the same noncommutative polynomials instantiated on matrices of any size. Another point of contact with our work is the consideration of relations between the sets in the sequence across dimensions, such as matrix-convex combinations which have been formalized and studied in this literature. Our notion of free descriptions is more general than the ones in this literature however, and it allows us to derive more flexible families of freely-described sets which are adapted to the structure underlying a broader range of applications. Further, the relations between sets in different dimensions we consider in this paper are less restrictive than matrix convexity, and they yield more general families of sets than free spectrahedra (free spectrahedra may be obtained in our framework via particular instantiations of description spaces, see Section 4.4).

### v3

@@ 1.1.3 Invariant Conic Programs
Our framework yields structural results not only for descriptions of convex sets but also for sequences of optimization problems over them. Specifically, we use representation stability to unify and generalize a number of previous results which show that various sequences of invariant conic programs indexed by dimension can be solved in constant time. Such sequences of programs arise in several applications including extremal combinatorics [3, 4] and quantum information [5, 6]. Concretely, the programs that arise in certifying homomorphism density inequalities over graphs are invariant under symmetric groups of increasing sizes that relabel the vertices of the graphs involved. Similarly, optimizing traces of matrix polynomials yields programs that are invariant under the unitary or orthogonal groups of increasing sizes that conjugate the matrices involved. We now recall symmetry reductions of invariant programs and explain how constant-sized reductions arise from representation stability.
@@
Let $\{\mathbb{U}_{n}\}$ be a sequence of $\{\mathsf{G}_{n}\}$ representations and $\{\mathcal{K}_{n}\subseteq\mathbb{U}_{n}\}$ a sequence of convex cones. For $t\in\mathbb{N}$, we say that the sequence $\{\mathcal{K}_{n}^{\mathsf{G}_{n}}\subseteq\mathbb{U}_{n}^{\mathsf{G}_{n}}\}$ admits a constant-sized description for $n\geq t$ if there exists a single finite-dimensional vector space $\mathbb{U}$ containing a cone $\mathcal{K}$, linear maps $T_{n}\colon\mathbb{U}\to\mathbb{U}_{n}^{\mathsf{G}_{n}}$, and subspaces $\mathbb{L}_{n}\subseteq\mathbb{U}$ such that $\mathcal{K}_{n}^{\mathsf{G}_{n}}=T_{n}(\mathcal{K}\cap\mathbb{L}_{n})$ for all $n\geq t$.
Note that Definition 1.10 does not require $\{\mathbb{U}_{n}\}$ to be a consistent sequence, nor in particular that $\mathbb{U}_{n}$ is embedded into $\mathbb{U}_{n+1}$. Proofs of constant-sized symmetry reductions in the literature have implicitly proceeded in a case-by-case manner by showing that the relevant cones, including symmetric PSD and relative entropy cones, have a constant-sized description in the sense of our Definition 1.10, see [25, 3, 26, 27]. In Section 5, we explain how these constant-sized descriptions can be generalized and derived systematically from an interplay between representation stability and the structure of the cones in question. To that end, we use a stronger form of representation stability known as uniform representation stability, which shows that the whole decomposition into irreducibles of the sequences of representations involved stabilize, rather than only their spaces of invariants; see Section 2.5. Our approach allows us to prove the existence of constant-sized descriptions for invariant sections of the following cones under the actions of the classical Weyl groups. The latter consist of the sequence of permutation groups $\{\mathsf{S}_{n}\}$, signed permutation groups $\{\mathsf{B}_{n}\}$, and even-signed permutation groups $\{\mathsf{D}_{n}\}$, see Table 1 for definitions.
Theorem 1.11 (Informal, see Theorem 5.1).
@@
The sequences of classical Weyl groups are the groups of permutations $\{\mathsf{S}_{n}\}$, signed permutations $\{\mathsf{B}_{n}\}$, or even signed permutations $\{\mathsf{D}_{n}\}$. Here $\mathrm{Sym}^{2}_{+}(\mathbb{V})$ denotes the collection of nonnegative quadratic forms on $\mathbb{V}$. For example, Theorem 1.11 shows that the sections of the PSD cones $\{\mathbb{S}^{n}_{+}\}$ invariant under the usual action of the classical Weyl groups by conjugation admit constant-sized descriptions. Theorem 1.11 is used to derive constant-sized descriptions for cones of invariant sums-of-squares in Theorem 1.12 below. In the following result, we consider polynomials in $\binom{n+k-1}{k}$ variables identified with $k$-multisets of $n$ letters, and show that the cones of sums of squares of such polynomials modulo any sequence of ideals admit constant-sized descriptions. The proof of the following result is given in Section 5.1.
@@
$\mathcal{I}_{n}\subseteq\bigoplus_{d\geq 0}\mathrm{Sym}^{d}\left(\mathrm{Sym}^{k}\mathbb{R}^{n}\right)\cong\mathbb{R}[x_{i_{1},\ldots,i_{k}}]_{1\leq i_{1}\leq\cdots\leq i_{k}\leq n}=:\mathbb{V}_{n}$
be $\mathsf{G}_{n}$-invariant ideals, and let $\mathbb{U}_{n}=\mathrm{Sym}^{\leq 2d}(\mathrm{Sym}^{k}\mathbb{R}^{n})/\mathcal{I}_{n}$. Consider the sums-of-squares cones $\mathrm{SOS}_{\mathbb{U}_{n}}=\{f\in\mathbb{U}_{n}:f\textrm{ is a sum of squares mod }\mathcal{I}_{n}\}$. Then the sequence $\{\mathrm{SOS}_{\mathbb{U}_{n}}^{\mathsf{G}_{n}}\}$ admits a constant-sized description for $n\geq 2kd$ if $\mathsf{G}_{n}=\mathsf{S}_{n}$ or $\mathsf{B}_{n}$, and for $n\geq 2kd+1$ if $\mathsf{G}_{n}=\mathsf{D}_{n}$.
Theorem 1.12 generalizes a number of results from the literature.
  • Setting $k=1$, $\mathcal{I}_{n}=(0)$, and $\mathsf{G}_{n}=\mathsf{S}_{n}$, we recover [25, Thms. 4.7, 4.10].
  • Setting $k=1$, $\mathcal{I}_{n}=(0)$, and $\mathsf{G}_{n}=\mathsf{B}_{n}$ or $\mathsf{D}_{n}$, we recover [26, Cor. 3.23].
  • For $k\geq 2$, setting $\mathcal{I}_{n}=(x_{I}-x_{I}^{2},x_{J})_{\begin{subarray}{c}I\in\binom{[n]}{k}\\ J\not\in\binom{[n]}{k}\end{subarray}}$ where $\binom{[n]}{k}$ denotes the set $\{(i_{1}<\ldots<i_{k}):i_{j}\in[n]\}$, and setting $\mathsf{G}_{n}=\mathsf{S}_{n}$, we recover [3, Thm. 2.4].
Theorem 1.12 generalizes all of these to include any of the classical Weyl groups and any sequence of invariant ideals.
Theorem 1.12 can be applied to derive constant-sized SDPs to certify graph homomorphism density inequalities [3, 4]. Many problems in extremal combinatorics can be recast as proving polynomial inequalities between homomorphism densities of graphs, which is the fraction of maps between the vertex sets of two graphs that define graph homomorphisms. A simple example is Mantel’s theorem, which states that the maximum number of edges in a triangle-free graph is $\lfloor n^{2}/4\rfloor$. Razborov proposed a method of certifying such inequalities using flag algebras [28], which were shown in [3, 4] to be sums-of-squares certificates of certain symmetric polynomial inequalities. Razborov’s flags are interpreted in [3, §3] as “free” spanning sets for spaces of symmetric polynomials. Formally, they are freely-described elements in the sense of Definition 1.4. Our framework shows that both the existence of such freely-described spanning sets and the resulting constant-sized SDPs are consequences of representation stability.
@@
1.1.4 Dimension-Free Convex Regression
As our final contribution, we apply our framework to develop an algorithm for a dimension-free analog of the convex regression problem. In the classic setting of this problem, the objective is to fit a convex function $f\colon\mathbb{V}\to\mathbb{R}$ to a dataset $\{(x_{i},y_{i})\}_{i=1}^{D}$ of inputs $x_{i}\in\mathbb{V}$ and outputs $y_{i}\in\mathbb{R}$ such that $f(x_{i})\approx y_{i}$. In our dimension-free extension of this problem, we have a sequence $\{\mathbb{V}_{n}\}$ of vector spaces, usually of growing dimension, and data $\{(x_{i},y_{i})\in\mathbb{V}_{n_{i}}\oplus\mathbb{R}\}$ in finitely-many dimensions $n_{i}$. We then seek an infinite sequence of convex functions $\{f_{n}\colon\mathbb{V}_{n}\to\mathbb{R}\}$ satisfying $f_{n_{i}}(x_{i})\approx y_{i}$ in the dimensions in which data is available and which, moreover, is freely-described.
To fit a freely-described sequence of functions, we begin by endowing the vector spaces $\{\mathbb{V}_{n}\}$ containing the training data with the structure of a consistent sequence based on the symmetries of the problem at hand and the relations between problem instances in different dimensions. We then select description spaces $\{\mathbb{W}_{n}\},\{\mathbb{U}_{n}\}$ and cones $\{\mathcal{K}_{n}\subseteq\mathbb{U}_{n}\}$ with respect to which our convex functions are to be described as in (ConicSeq), with freely-described vectors and linear maps parametrizing the desired family of freely-described functions; richer description spaces generally yield more expressive families of convex functions, but fitting a function from such a family is more expensive and requires data in higher dimensions by Theorem 3.5.
@@ 1.1.4 Dimension-Free Convex Regression
The $\ell_{\pi}$ norm of a vector is defined for vectors of any length, is invariant under signed permutations, remains unchanged under zero-padding of its input, and can only increase if we append any other entry to a given input. It is therefore both intersection and projection compatible with respect to the embeddings and projections of Example 1.5. The function (2) is well-defined for (positive-semidefinite) inputs of all sizes, is invariant under conjugation by orthogonal matrices, and its value is unchanged if we zero-pad its input. Thus, it is intersection-compatible on $\{\mathbb{S}^{n}\}$ with embeddings given by zero-padding. Given evaluations of these two functions on low-dimensional inputs, we use our procedure to fit semidefinite approximations to them by choosing sequences of PSD cones for our descriptions, and fitting a freely-described and compatible sequence of convex functions. Thus, we ensure that our approximations are also group-invariant and satisfy the same compatibility properties as the underlying ground-truth functions. Concretely, we fit an approximation to the $\ell_{\pi}$ norm using its values on 50 random vectors of length at most 2, and fit an approximation to (2) using its values on 200 random PSD matrices of size at most $4\times 4$. To evaluate the quality of our approximations, we further generate $10^{3}$ random points in dimensions up to 20 and compare the value of our approximation to the ground truth functions on these points. We detail our description spaces, our fitting algorithm, and our data generation process in Section 6.3. Figure 1 shows the error in our semidefinite approximations for different dimensions when we search over all freely-described sets in our family when fitting to data, and when we search only over the smaller family of compatible sequences. We see that imposing compatibility ensures that the error increases gracefully when extending to higher dimensions.
@@
Obtaining semidefinite approximations for quantum information theoretic quantities such as (2) facilitates the use of semidefinite programming solvers in convex optimization problems involving quantum entropy [32, 33, 34, 35]. We mention that the authors of [36] derived semidefinite approximations of the quantum entropy and related functions from an analytic perspective. We show in Section 4.5 that their description is dimension-free but does not certify compatibility, in contrast to the description we obtain from data using the preceding procedure.
@@ 1.2 Related Work
There is a large literature as well as a systematic framework on extended formulations in which conic descriptions of a convex set in a fixed dimension are investigated; see [37] for a review. The goal in this body of work is to express ‘complicated’ convex sets in $\mathbb{R}^{d}$ (e.g., polyhedra with many facets) as linear images of affine sections of ‘simple’ convex cones in a space that is not much larger than $d$. This framework is also applied to study equivariant lifts of group-invariant convex sets, which are descriptions of the form (Conic) consisting of group-invariant cones, vectors, and linear maps, i.e., conic descriptions that make evident the group invariance of the convex set; see [38] and [37, §4.3]. These are precisely the type of descriptions we consider for each of the convex sets in our sequences. Moreover, while the literature on extended formulations is typically articulated in the setting of a convex set in a fixed dimension, many results in the area implicitly concern descriptions of a sequence $\{\mathcal{C}_{n}\}$ of convex sets and the complexity of these descriptions as a function of $n$ [39, 38, 40, 41]. Thus, there are several points of contact with the present paper. In fact, many (though not all) descriptions proposed in the literature on extended formulations can in fact be instantiated in any dimension, and are moreover dimension-free in our sense as we show in Section 4. More broadly, to the best of our knowledge, descriptions of convex sets that arise from a systematic consideration of relations between dimensions have not been studied previously. By bringing such considerations to the fore, the present effort elucidates the representation-theoretic phenomena and their interaction with the convex-geometric aspects underpinning convex sets that can be instantiated in any dimension.
@@
A broad research program pursued in several areas involves the study of “matrix” or noncommutative analogues of classical “scalar” or commutative objects. Examples include random matrix theory and free probability [42] in which matrix-valued random variables and their limits are the object of study as opposed to scalar-valued ones; and noncommutative algebraic geometry [13], in which polynomials in noncommuting variables and their evaluations on matrices are studied as opposed to standard polynomials in commuting variables that are evaluated on scalars [43, 44, 45, 46]. Applying this program to convex sets yields matrix-convex sets and free spectrahedra as in Example 1.2(b). We refer the reader to [13, 14] for surveys and [47, 48] for some applications. In analogy to the present paper, results in this area explicitly pertain to sequences of sets which are “freely-described”, in the sense that their description can be instantiated in any dimension. For example, free spectrahedra are sequences of sets described by a single linear matrix inequality, and free algebraic varieties are defined by the same noncommutative polynomials instantiated on matrices of any size. Another point of contact with our work is the consideration of relations between the sets in the sequence across dimensions, such as matrix-convex combinations which have been formalized and studied in this literature. Our notion of dimension-free descriptions is more general than the ones in this literature however, and it allows us to derive more flexible families of freely-described sets which are adapted to the structure underlying a broader range of applications. Further, the relations between sets in different dimensions we consider in this paper are less restrictive than matrix convexity, and they yield more general families of sets than free spectrahedra (free spectrahedra may be obtained in our framework via particular instantiations of description spaces, see Section 4.4).